2. Simplify the following expressions and write your answers in closed form.
Answer #1

21 2 Ei2 + 3 si izl izl Now Siz 여 ninti) (mell sw ninti) 2 2 arba nint) (zuel) + 3 n(n+1) & 3 12 n(n+1) n(n+1) [mei + 370 mel

3 WE i 2- 4 = (i3_4n if the above expression is и 3 (nentu) – + CM n2 (n+112 - ulnl n (n (+ y)2-4) 4 = n (n(n+1)2-16) si3-4 i

A thumbs up is appreciated.